find the sum of all positive integral values of a for which every solution to the system of equation x +ay=3 and ax+4y=6 satisfy the inequalities x>1,y>0

Grade:12

1 Answers

Sachin
5 years ago
  1. ax+a2y=3a
  2. ax + 4y=6
(1)-(2)=a2-4y=3a-6
a2y-3aa4y+6=0
For real D>=0
9+(4y(4y-6)>=0
Y>=3/4. 
Similarly find X 
And then sum all satisfied solutions